OpenStudy (anonymous):

which is an equation of the given line in standard form? (1/2,3) and (-4,-1) a.)-8x+9y=23 b.)-8x+9y=-23 c.)-8x+7y=25 d.)-9x+8y=-23

OpenStudy (blurbendy):

The first step is to find the slope: slope = y2 - y1 / x2 - x1 where y2 = -1 y1 = 1 x2 = -4 x1 = 1/2 Can you find the slope?

OpenStudy (anonymous):

kinda im that not the best at it

OpenStudy (blurbendy):

It's okay, show me what you got.

OpenStudy (anonymous):

ok so I know that you have to y1 -y2 and x1 -x2 so for y is 1/2 - -4= 5 and the x1 - x2 is 3-(-1) = 4 and that is what I got so far

OpenStudy (blurbendy):

-1 -3 = -4 -4-1/2 = -4.5 -4/-4.5 = 8/9 now put this in the form: y - y1 = 8/9(x - x1) y +1 = 8/9x + 32/9 y = 8/9x + 23/9 standard form = Ax + By = C multiply everything by 9 9y = 8x + 23 -8x +9y = 23